40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Mon Jun 11 09:26:37 2018, Disha Kapasi, Update, General, Current Driver v2 and Frequency noise of the beat note Current_Driver_Circuit_-_v2.pdfnewbeatnote.pdf
    Reply  Tue Aug 14 23:21:01 2018, johannes, Update, General, Current Driver v2 20180814_160659.jpgn_channel_driver.PNGp_channel_driver.PNGvis_mods.pdf
       Reply  Wed Aug 15 23:34:38 2018, johannes, Update, General, Cooling down 
          Reply  Mon Aug 20 02:35:06 2018, johannes, Update, General, We're cool 
Message ID: 2139     Entry time: Mon Aug 20 02:35:06 2018     In reply to: 2134
Author: johannes 
Type: Update 
Category: General 
Subject: We're cool 

We're cool. For a couple days now. The CTC100 was set to maintain the thermal shields of both cavities at 120K.

With the current drivers in limbo I couldn't check what worried me most: That the mirrors may have plopped off due to insufficient heat transfer. I brought both drivers back to working order, one remains not installed for testing, but I installed the other one and fired up the lasers one after the other to look for back-reflections.

Upon turning the light on I saw reflected light immediately. A bit misaligned but surprinsingly close to the incident beam. The following procedure proved quite successful for both paths, it took less than 10 minutes each to recover the (frontal) alignment:

  1. From before I knew that I had to move the incident beam way up, which was done in a controlled way to not lose the back-reflection
  2. Started looking for dips in the DC reflected light and quickly found them in both cases (which means all mirrors are still attached, the point of this test yes)
  3. Found a somewhat strong dip, place a fast lens at the output viewport and positioned the CCD camera while dithering the laser frequency to capture the transmitted beam on the sensor
  4. The mode shape hints at the misalignment. In this case, both were still mostly off in pitch
  5. Corrected alignment and checked for visible lower order modes; rinse and repeat until 00 is identified
  6. I didn't align the transmitted path yet, instead I dithered the laser frequency again and maximized the dip in the DC reflection
  7. Did a quick test lock for both. Both paths lock just fine, transmission doesn't show any signs of clipping and positions on the front viewports are a bit high but in the clear. Because of the tubes of the acoustic enclosure it's a bit hard to say the same thing for transmission though.

It's a major relief to know that the mirrors stayed on and that the bonding strategy worked as intended.

Next, the transmission path will need to be realigned. Last week we received the 4" extruded M2ISS PD mounts, the PCBs, and the circuit parts. Anchal and I will begin assembly and if all goes well we will have shot noise limited TransPD readout soon.

Design of the AC-coupled ISS servo also moves forward and we'll probably be ready to order the parts next week.

ELOG V3.1.3-